A 1.00 L solution contains 2.00*10^-4M Cu(NO3)2 and 1.20*10^-3M ethylenediamine (en). The Kf for Cu(en)2 2+ is 1*10^20 . What is the concentration of Cu2+ (aq) in the solution.
Answer #1

The formation constant for   Cu(en)22+ is 1*1020
                 Cu2+   +    2 en ----------->   Cu(en)22+

I (M) :   2.00*10-4    1.6*10-3                        0

C(M) :        -x                -2x                          +x

           --------------------------------------------------

E(M) 2.00*10-4 - x   1.6*10-3 - 2x                 x

            Kf = [Cu(en)22+] / [ Cu2+][en]2

     1*1020 = x (1.6*10-3 - 2x) / (2.00*10-4 - x)

                 x = 5.00*1019 M